OpenStudy (anonymous):

A substance is found to have a melting point of 80°C. The vapour pressure of the liquid is 40 Torr at 108.2°C and 60 Torr at 119°C. The vapour pressure of the solid is 1 Torr at 52.6°C. Using only this information, calculate

OpenStudy (anonymous):

a)the boiling point of the substance and its enthalpy and entropy of vaporization at the boiling point b)the vapour pressure at the melting point; c) the enthalpy of sublimation and fusion of the solid, assuming that the melting and triple point temperatures are the same

OpenStudy (anonymous):

A little more work on this one, couldn't quite get c)

OpenStudy (jadedry):

Hmm, this definitely looks trickier. Give me a few minutes!

OpenStudy (anonymous):

Alright i'll be here attempting to wrap my head around c)

OpenStudy (jadedry):

Alright i'm going to focus on c with you, the first two seem okay, you just need to rearrange the equation we used last time as both of these are talking about vaporization.

OpenStudy (jadedry):

I'm looking this up and I came up with a helpful tidbit: To estimate sublimation use the vapor pressures of the solid phase, (we only have 1? I suppose the melting point of 80 degrees is a hint to find a second value. Also, brb something came up.

OpenStudy (jadedry):

to estimate vaporization use the vapor pressures of the liquid phase. (have two of those, that's good.) and fusion = sublimation - vaporization

OpenStudy (anonymous):

oh alright that's a lot of helpful tidbits

OpenStudy (jadedry):

I think the main problem is finding sublimation, we can get vaporization and hence fusion after that.

OpenStudy (jadedry):

Because we just input the given values in the P1/2 and T1/2 variables. I suppose we could consider using 80 degrees or a bit less, just because it's melting doesn't mean it's not a solid. It must be important otherwise they would not have given us that info.

OpenStudy (jadedry):

Wait, what am I saying. Ignore my previous statement let me see...

OpenStudy (anonymous):

Aha alright, I'm reading what you're writing I guess I'll forget it all

OpenStudy (jadedry):

if we use the equation using the solid value. \[Ln (760/760) =- \frac{ H }{ R } (\frac{ 1 }{ 325.75 } - \frac{ 1 }{ 353 })\] (because it has a vapor pressure of 1 at 52.6°c) So we solve the equation for H and we get vaporization!

OpenStudy (jadedry):

I think I was on the right track? but maybe not explaining myself in the right way.

OpenStudy (jadedry):

1 torr, lmao i'm so tired. it would be 1/760 , not 1/1

OpenStudy (anonymous):

it's alright I understood what you meant. I'm starting to understand it better now (y) , thanks alot

OpenStudy (jadedry):

Good! happy to help. c: I should go back to my studies, if you want any help just give me a ping and I'll reply when I can. ;u;

OpenStudy (anonymous):

Alright good luck my man, I'll probably pass out soon it's pretty late here

OpenStudy (jadedry):

Alright, cheers, and good luck with the rest of your studies!